OpenStudy (masierr):

The equation for the magnitude M of an earthquake is M = log (I/K) where I is the intensity of the earthquake and K is a constant. The magnitude of a 1995 earthquake was 9.0 and the magnitude of a 2004 earthquake was 7.0. Which of the following compares (I)1995, the intensity of the 1995 earthquake with (I) 2004, the intensity of the 2004 earthquake? Multiple choice answers. A. (I)1995= 100(I) 2004 B. (I)1995 = 2(I)2004 C. (I)1995 = (I)2004 + 2 D. (I) 1995 = (I)2004 + 100 E. (I)1995 = 2(I)2004 + 100

OpenStudy (g3n3sis08):

I think its C. M = log (i/k) M = log i - log k a. 9 = 1995i - log k b. 7= 2004i - log k So a-b to compare difference therefore, 9-7 = 1995i - log k - (2004i - log k) 2 = 1995i - log k - 2004i +log k " log k gets cancelled" so, 2 = 1995i - 2004i Ans.. 2 + 2004i = 1995i

OpenStudy (mathmate):

We are given M(I)=log(I/K)=log(I)-log(K) where K is a constant. We also know that \(M_{1995}=9=log(I_{1995})-log(K)\), and \(M_{2004}=7=log(I_{2004})-log(K)\) Subtract the two gives \(M_{1995}-M_{2004}=9-7=2=log(I_{1995})-log(I_{2004})=log((I_{1995}/I_{2004})\) or \(log((I_{1995}/I_{2004})=2\) Assuming given equation uses log to base 10, raise to power 10 on both sides give \(I_{1995}/I_{2004}=10^2\) and the result follows.